OpenStudy (anonymous):

**WILL MEDAL** A line has a slope of 9/4 and a y-intercept of -14. Which equations represent the line? Choose all answers that are correct. A. -9x+4y=-56 B. 9x+4y=56 C. y=9/4x-14 D. y=9/4x+14

OpenStudy (anonymous):

C.

OpenStudy (anonymous):

C is correct

OpenStudy (dan815):

i give up -.-

OpenStudy (anonymous):

Thx

OpenStudy (dan815):

forget learning

OpenStudy (anonymous):

np

OpenStudy (dan815):

come lets all just cheat u got more questions? ill do your whole test/assignment for u

OpenStudy (anonymous):

Yea sure ig but this one is suppose to have 2 answers

OpenStudy (dan815):

ya okay

OpenStudy (dan815):

thought abut rearranging c?

OpenStudy (dan815):

how about subtracting 9/4x from both sides and multiplying both side by 4,

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

Sorry u confused me

OpenStudy (dan815):

okay.. look at c)

OpenStudy (dan815):

y=9/4 x - 14

OpenStudy (anonymous):

Ok

OpenStudy (dan815):

both sides of the equations are equal, you subtract 9/4x from both sides and it wont change anything right..

OpenStudy (dan815):

like if A=B then A-2=B-2 is still true

OpenStudy (dan815):

y -9/4x=9/4 x - 14 -9/4x

OpenStudy (dan815):

now the right side simplifies

OpenStudy (dan815):

y -9/4x=- 14

OpenStudy (dan815):

what happens now if u multiply both sides by 4?

OpenStudy (anonymous):

Sorry computer froze

OpenStudy (anonymous):

y-37.6=-56?

OpenStudy (anonymous):

@dan815 ?